# Are AI Detectors Good Enough? A Survey on Quality of Datasets With Machine-Generated Texts

### v1

@@ 5 Experiments
Dataset DeBERTa Binoculars DetectGPT
@@
M4 0.987 0.171 0.381
SemEval24 Mono 0.999 0.943 0.983
SemEval24 Multi 0.997
@@
Table 3: Classification results with different detectors estimated using $F_{1}$-score. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
From each dataset, we sampled 1000 documents, balanced between two classes. As we dealt with multilingual datasets, we chose to finetune a multilingual model supporting all the presented languages, namely mDeBERTa-v3-base. Information about training can be found in the Appendix D.To evaluate the quality of baselines, Binoculars and Fast-DetectGPT, we launched falcon-rw-1b Almazrouei et al. (2023) and gpt-neo-2.7B Black et al. (2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English. Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used RoBERTa-base, just as the authors of original paper. In the experiment Perturbations and Shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages Wang et al. (2024a). Finally, in the experiments with attention maps we used LLaMA2-7B to obtain attention scores. We took 3 heads from Layer 15 and averaged them, as heads on this layer showed very distinct attention patterns.
6 Results
Dataset $\text{KL}_{\text{TTS}}$downarrow Attention Columns (h / m) $\Delta_{\text{shift}}\downarrow$ $\text{KL}_{\text{shuffle}}\downarrow$
GPT-2 0.014 3.430 / 4.094 0.084 1.255
HC3 0.053 0.459 / 0.967 0.264 1.167
GhostBuster 0.053 2.822 / 2.988 0.024 0.359
MGTBench 0.043 1.961 / 2.639 0.031 0.421
MAGE 0.011 2.289 / 2.166 0.094 0.310
M4 0.036 3.842 / 2.256 0.107 0.483
SemEval24 Mono 0.012 1.540 / 0.766 0.191 2.576
SemEval24 Multi 0.001 2.123 / 0.830 0.059 2.046
RuATD 0.007 1.631 / 1.391 0.315 14.028
DAGPap22 0.083 0.637 / 0.675 0.039 0.472
PAN24 0.053 3.463 / 2.588 0.050 0.331
AuTex23-en 0.021 3.179 / 2.740 0.110 4.331
AuTex23-es 0.001 3.072 / 3.244 0.105 1.306
IberAuTex 0.012 2.049 / 1.946 0.223 5.516
MGT-1 Mono 0.019 2.070 / 1.783 0.031 0.587
MGT-1 Multi 0.006 3.313 / 3.117 0.027 0.522
Table 4: Calculated statistics on texts from chosen datasets. In “Attention Columns” we show the mean difference between the highest attention column and the second-placed. The first value is for human texts; the second value is for machine-generated. Some values for $\text{KL}_{\text{TTS}}$ are underlined, because texts are too short, see Section 7.
Figure 2: Topological Time Series for different datasets. These datasets obtained highest $\text{KL}_{\text{TTS}}$. The results for the remaining datasets selected in this paper can be found in Figure 4.
Figure 3: Attention maps on two excerpts from DAGPap22, Layer 15, Head 15.
@@ 6 Results
The results of comparison of the designed features on the test parts of the selected datasets are presented in Table 4. Regarding the TTS score, previous works have shown that texts of different origin have different PHD values, however this result was obtained for GPT-2 model, meanwhile current advanced models generate texts of much better quality, therefore PHD values become more similar. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7.
@@
The second column of Table 4 is based on attention map analyses. Here, on the left are the mean difference values of the human excerpts, on the right are the generated ones. When the attention maps of texts of different origin show different patterns, it also becomes easier for detectors to separate the texts. In the remaining columns we list the statistics observed on modified texts, and for both of these the lower the better, as this reflects the similar degree of resilience of the generated and human texts to adversarial attacks. Qualitatively generated data with no bias should take values close to human.
@@
Finally, in Table 3 we show the results of applying modern classifiers to the chosen test parts. For instance, on the datasets that had low values in Table 4, a quality close to 1 can be achieved, which indicates the clear presence of detector bias towards them, or a structural feature that is too obvious for the model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables we can estimate which set has better quality data and which has lower quality data.
@@ 7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with high value of it. While GhostBuster and PAN24 received such high score due to discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es, as they do not fit the criteria, see Table 1. On top of that, it has already been shown that texts must be of sufficient length Gritsay et al. (2022) to build reliable detectors.
Regarding attention maps, in Fig. 3, we depict an example of distinct patterns where human text has only one separate attention column besides attention on the first token, while machine-generated text has multiple attention columns.
Analysing the values in the table 4, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools for evaluating data quality, e.g. Zipf’s law Powers (1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. Also, datasets that collect machine-generated content may provide utility for the more general two other purposes as well. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
@@ 9 Limitations
Appendix A Additional Results on Topological Time Series
In Figure 4, it can be seen the plotted Topological Time Series distributions for all the selected datasets for the currect research.
Figure 4: Topological Time Series for remaining datasets from current review. We omitted the results for AuTex23en, because virtually all texts there had the dimension of 0.
Appendix B Data Description
More detailed description with information on sources, topics and years of the datasets selected in this paper from competitions in Table 5 and research articles presented in Table  6.
@@ Appendix B Data Description
GPT2 2019 WebText GPT-2-117M, GPT-2-345M, GPT-2-762M, GPT-2-1542M
HC3 2023 ELI5, WikiQA, Wikipedia, Medicine, Finance ChatGPT (gpt-turbo-3.5)
GhostBuster 2023 Student Essays, News Articles, Creative Writing ChatGPT (gpt-3.5-turbo), Claude
MGTBench 2024 Student Essays, News Articles, Creative Writing ChatGLM, Dolly, ChatGPT-turbo, GPT4All, StableLM, Claude
MAGE 2024 Opinions, Reviews, News, QA, Story Generation, Commonsense Reasoning, Knowledge Illustration, Scientific Writing ChatGPT (text-davinci-002, text-davinci-003, gpt-turbo-3.5), LLaMA, GLM-130B, FLAN-T5, OPT, BigScience, EleutherAI
M4 2024 Wikipedia, Reddit ELI5, WikiHow, PeerRead, arXiv abstract ChatGPT (text-davinci-003, gpt-turbo-3.5), Cohere, Dolly-v2, BLOOMz
Table 6: More detailed descriptive statistics about domains and generators of the chosen datasets from papers.
Appendix C Evaluation results of competitions
Table 7 shows the winning scorers in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of metrics on the RuATD dataset, which can be explained by the limited number of available high-quality language models in Russian during the competition.
@@ Appendix C Evaluation results of competitions
AuTex-en Macro-F1 0.809
AuTex-es Macro-F1 0.708
@@
Appendix D Hyperparameters
@@
The training was conducted on NVIDIA GeForce RTX 3090. See hyperparameters in Table 8.

### v2

@@ 5 Experiments
From each dataset, we sampled 1000 documents from the test set, balanced between two classes. Regarding baselines, we fine-tuned mdeberta-v3-base for each dataset and evaluated the model, information about training can be found in the Appendix C. To evaluate the quality of baselines, Binoculars and Fast-DetectGPT, we utilised falcon-rw-1b Almazrouei et al. (2023) and gpt-neo-2.7B Black et al. (2021) respectively. It is worth noting that with the last two methods we were only able to measure quality for samples in English.
Our objective was to show that datasets of lower quality have shifts that will be easily recognised by the models ”from the first step”, hence we have not performed any hyperparameter tuning, only one iteration of fine-tuning and testing of the underlying models. In the experiment with topological features we used roberta-base, just as the authors of the original paper. In the experiment with perturbations and shuffling, the multilingual-e5-large encoder was used to build embeddings of texts, which shows high metrics on encoding high-resource languages Wang et al. (2024a).
6 Results
Dataset $\text{KL}_{\text{TTS}}$downarrow $\text{PHD}_{\text{human}}$ $\text{PHD}_{\text{machine}}$ $\Delta_{\text{shift}}$downarrow $\text{KL}_{\text{shuffle}}$downarrow
GPT-2 0.014 9.23 $\pm$ 1.98 10.27 $\pm$ 1.84 0.084 1.255
HC3 0.053 8.76 $\pm$ 1.83 7.38 $\pm$ 1.05 0.264 1.167
GhostBuster 0.053 9.84 $\pm$ 1.18 9.76 $\pm$ 1.15 0.024 0.359
MGTBench 0.043 8.77 $\pm$ 1.31 9.97 $\pm$ 1.02 0.031 0.421
MAGE 0.011 9.8 $\pm$ 2.14 9.38 $\pm$ 3.04 0.094 0.310
M4 0.036 7.26 $\pm$ 1.99 8.59 $\pm$ 1.4 0.107 0.483
OutFox 0.025 8.96 $\pm$ 1.21 11.48 $\pm$ 1.13 0.095 0.237
TweepFake - 9.02 $\pm$ 3.19 8.12 $\pm$ 4.02 0.116 1.001
SemEval24 Mono 0.012 9.11 $\pm$ 1.19 9.41 $\pm$ 1.2 0.191 2.576
SemEval24 Multi 0.001 9.65 $\pm$ 1.81 9.42 $\pm$ 1.44 0.059 2.046
RuATD 0.007 7.33 $\pm$ 1.4 7.46 $\pm$ 1.41 0.315 14.028
DAGPap22 0.083 8.35 $\pm$ 1.33 7.48 $\pm$ 2.01 0.039 0.472
PAN24 0.053 9.4 $\pm$ 1.05 8.52 $\pm$ 1.59 0.050 0.331
AuTex23 Eng 0.021 8.07 $\pm$ 2.26 8.1 $\pm$ 2.68 0.110 4.331
AuTex23 Esp 0.001 9.16 $\pm$ 3.49 9.25 $\pm$ 3.26 0.105 1.306
IberAuTex 0.012 9.33 $\pm$ 2.45 8.47 $\pm$ 2.73 0.223 5.516
MGT-1 Mono 0.019 9.19 $\pm$ 1.75 8.96 $\pm$ 2.24 0.031 0.587
MGT-1 Multi 0.006 8.76 $\pm$ 1.85 8.6 $\pm$ 2.29 0.027 0.522
Table 2: Calculated statistics on texts from chosen datasets. Some values for $\text{KL}_{\text{TTS}}$ are underlined, because texts are too short, see Section 7 and TTS for almost all texts in TweepFake is equal to 0.
Dataset BERT Binoculars DetectGPT
@@ 6 Results
M4 0.987 0.871 0.881
OutFox 0.901 0.692 0.707
SemEval Mono 0.991 0.913 0.924
SemEval Multi 0.994
@@
Table 3: Classification results with different detectors estimated using $F_{1}$-score. BERT here is DeBERTa-based classifier. Binoculars and DetectGPT work only with English texts, thus we could not apply them to datasets with non-English texts.
@@
The results of the comparison of the designed features on the selected datasets are presented in Table 2. Regarding the PHD and TTS score, in previous works it was shown that texts from language models have smaller PHD values than human-written ones, however this result was obtained for GPT-2, GPT-3.5 and OPT models, and this trend could change for more recent language models, that generate more human-resembling texts. If texts of different origin have high $\text{KL}_{\text{TTS}}$, it means that it is easier for a detector to separate such texts. $\text{KL}_{\text{TTS}}$ is also constrained for shorter texts, see Section 7. As for PHD, we hypothesize that generated texts of good quality should have similar PHD with human-written ones. Additionally, we compare the distributions of PHD for all datasets, see Figure 3. Again, the distributions for texts of both origin should be similar, which holds for texts from SemEval, PAN24 and MGT-1.
@@
In the next columns, we list the statistics observed on modified texts, and for both of these the lower the better, as this reflects the similar degree of resilience of the generated and human texts to adversarial attacks. Qualitatively generated data with no bias should take values close to human.
@@
Finally, in Table 3 we show the results of applying modern classifiers to the chosen test parts. For instance, on the datasets with low values in Table 2, a quality close to 1 can be achieved, which indicates the clear presence of detector bias towards them, or a structural feature that is too obvious for the model. It is not possible to judge the quality of the data only by achieving $F_{1}$ values close to 1, but by combining the values of the two tables, we can estimate which set has better quality data and which has lower quality data.
@@ 7 Discussion
Regarding $\text{KL}_{\text{TTS}}$, on Fig. 2 we show 4 datasets with high value of it. While GhostBuster and PAN24 received such high score due to discrepancy on texts with higher dimensions, MGTBench and DAGPap22 did it due to the difference in distributions themselves. Note also that $\text{KL}_{\text{TTS}}$ may not perform well with very short texts, since the internal method of computing PHD requires sufficiently long texts for stable computation. Therefore, we discard $\text{KL}_{\text{TTS}}$ on RuATD and AuTex23-es and TweepFake, as they do not fit the criteria, see Table 1. On top of that, it has already been shown that texts must be of sufficient length Gritsay et al. (2022) to build reliable detectors.
Analysing the values in the Table 2, we can trace the presence of sufficiently high quality data in the selected datasets. The developed attributes in aggregate are able to reflect the quality of the generated dataset from different perspectives and angles. We propose to utilise these attributes in combination with other statistical tools for evaluating data quality, e.g., Zipf’s law Powers (1998).
Presented statistics can be utilised to estimate the quality of collections and to improve them. Also, datasets that collect machine-generated content may provide utility for the two more general purposes as well. First, high-quality generated data can be utilised to evaluate the quality of the causal model during training, as one of the training objectives to improve model answers and make it more human-like. Secondly, good detectors can help to clean training sets, as large proportion of low-quality generated texts in those sets can result in emerging biases towards incorrect structure and rubbish fragments in the output of the model in the future.
@@ 9 Limitations
Figure 3: PHD values on all datasets, except TweepFake and AuTex23 Spanish, texts from which were too short for proper calculation of PHD.
Appendix A Data Description
More detailed description with information on sources, topics and years of the datasets selected in this paper from competitions and research papers in Table 4
@@ Appendix A Data Description
Research papers datasets
GPT2 2019 WebText GPT-2
TweepFake 2019 Tweets Markov Chains, RNN, LSTM, GPT-2
HC3 2023 ELI5, WikiQA, Wikipedia, Medicine, Finance ChatGPT
GhostBuster 2023 Student Essays, News Articles, Creative Writing ChatGPT, Claude
MGTBench 2024 Student Essays, News Articles, Creative Writing ChatGPT. ChatGLM, Dolly, GPT4All, StableLM, Claude
MAGE 2024 Opinions, Reviews, News, QA, Story Generation, Commonsense Reasoning, Knowledge Illustration, Scientific Writing text-davinci-002, GPT-3.5, ChatGPT, LLaMA, GLM-130B, FLAN-T5, OPT, BLOOM, GPT-J-6B, GPT-NeoX-20B
M4 2024 Wikipedia, Reddit ELI5, WikiHow, PeerRead, arXiv abstract GPT-3.5, ChatGPT, Cohere, Dolly-v2, BLOOM
OutFox 2024 Student Essays ChatGPT, GPT-3.5, FLAN-T5
Shared tasks datasets
RuATD 2022 News, Social media, Wikipedia, Strategic Documents, Diaries M-BART, M2M-100, OPUS-MT, mT5, ruGPT2, ruGPT3, ruT5-Base
DAGPap 2022 Scopus papers Longformer Encoder-Decoder, GPT-3, Spinbot, GPT-Neo
AuTex 2023 Legal documents, Social media, How-to articles BLOOM, GPT-3, GPT-3.5
IberAuTex 2024 News, Reviews, Emails, Essays, Dialogues, Wikipedia, Wikihow, Tweets GPT-2, LLaMA, Mistral, Cohere, Claude, MPT, Falcon
PAN 2024 News Alpaca, BLOOM, Gemini, ChatGPT, gpt-4-turbo, LLaMA-2, Mistral, Qwen1.5, GPT-2
SemEval Mono 2024 Wikipedia, WikiHow, Reddit, arXiv, PeerRead, Student Essays GPT-3.5, GPT-4, Cohere, Dolly-v2, BLOOMz
SemEval Multi 2024 Wikipedia, WikiHow, Reddit, arXiv, PeerRead, Student Essays, News ChatGPT, GPT-3.5, GPT-4, LLaMA2, Cohere, Dolly-v2, BLOOM, Jais
MGT-1 Mono 2025 CNN, DialogSum, Wikipedia, WikiHow, Eli5, Finance, XSum, PubMed, SQuAD, IMDb, Reddit, arXiv, PeerRead text-davinci-002, GPT-3.5, ChatGPT, OPT, LLaMA3, BLOOM, FLAN-T5, Cohere, Dolly, Gemma, Mixtral
MGT-1 Multi 2025 CNN, DialogSum, Baike, WikiQA, WikiHow, Eli5, Finance, Psychology, XSum, PubMed, SQuAD, IMDb, Reddit, arXiv, PeerRead text-davinci-002, GPT-3.5, ChatGPT, gpt4o, GLM, GPT-J, GPT-Neo, OPT, LLaMA2, LLaMA3, BLOOM, FLAN-T5, Cohere, Dolly, Gemma, Mixtral, Jais
Table 4: More detailed descriptive statistics about domains and generators of the chosen datasets from competitions and research papers. ChatGPT is gpt-3.5-turbo, GPT-3.5 is text-davinci-003.
Appendix B Evaluation results of competitions
Table 5 shows the winning scores in the competitions reviewed in this paper. In the AuTex and IberAuTex competitions it was forbidden to use additional data to fine-tune the detection algorithms. In the other collections it was allowed, we can notice a high quality near perfect in them. We should note the low value of metrics on the RuATD dataset, which can be explained by the limited number of available high-quality language models in Russian during the competition.
@@ Appendix B Evaluation results of competitions
AuTex23-en Macro-F1 0.809
AuTex23-es Macro-F1 0.708
@@
Figure 4: Topological Time Series on all datasets.
Appendix C Hyperparameters
@@
The training was carried out on NVIDIA GeForce RTX 3090. See hyperparameters in Table 6.